# ORBIT-2: Scaling Exascale Vision Foundation Models for Weather and Climate Downscaling

### v1

@@
Abstract
Sparse observations and coarse-resolution climate models limit effective regional decision-making, underscoring the need for robust downscaling. However, existing AI methods struggle with generalization across variables and geographies and are constrained by the quadratic complexity of Vision Transformer (ViT) self-attention. We introduce ORBIT-2, a scalable foundation model for global, hyper-resolution climate downscaling. ORBIT-2 incorporates two key innovations: (1) Residual Slim ViT (Reslim), a lightweight architecture with residual learning and Bayesian regularization for efficient, robust prediction; and (2) TILES, a tile-wise sequence scaling algorithm that reduces self-attention complexity from quadratic to linear, enabling long-sequence processing and massive parallelism. ORBIT-2 scales to 10 billion parameters across 32,768 GPUs, achieving up to 1.8 ExaFLOPS sustained throughput and 92–98% strong scaling efficiency. It supports downscaling to 0.9 km global resolution and processes sequences up to 4.2 billion tokens. On 7 km resolution benchmarks, ORBIT-2 achieves high accuracy with $R^{2}$ scores in range of 0.98–0.99 against observation data.
I Problem Overview
@@
Downscaling bridges this gap by translating coarse resolution global climate model outputs into fine-scale outputs [1]. This process is critical across a wide range of sectors, including agriculture [2], water resources [3], infrastructure and urban planning [4], energy systems [5], and extreme event forecasting [6]. Despite its importance, downscaling remains both scientifically and computationally challenging. It requires physically consistent predictions from massive, high-dimensional spatiotemporal data while maintaining accuracy across diverse regions. These challenges highlight the urgent need for scalable, high-fidelity downscaling approaches to enable effective climate services, disaster preparedness, and policy planning.
@@
Traditional downscaling methods fall into two main categories: dynamical [7], which uses nested physical models to simulate fine-scale processes but is computationally intensive and limited to regional domains; and statistical [8, 9], which is computationally efficient but often lacks physical fidelity and generalizability. More recently, artificial intelligence (AI) has emerged as a powerful alternative, offering the potential for high-resolution predictions with significantly lower inference costs [10, 11, 12]. Task-specific deep learning models have been developed to learn mappings between coarse and fine-resolution fields, typically falling into two categories: non-generative models [13], which are efficient and stable but may lack fine spatial detail, and generative models, such as diffusion-based approaches [14, 15, 16], which produce sharper outputs at the expense of much higher computational and training costs. Despite promising results, most task-specific models must be retrained for each variable, resolution, or region and often struggle to generalize across diverse, physically distinct climate variables [17].
@@
Figure 1: A generalized AI architecture diagram for state-of-the-art downscaling foundation models. Note that upsampling is used for each channel prior to training blocks to reduce downscaling uncertainty.
@@
To overcome the limitations of task-specific models, recent efforts have introduced foundation models such as Prithvi [18] and ClimateLearn [19], which employ multi-task Vision Transformer (ViT)-based architectures to support downscaling across variables and geographic regions. While these models represent an important step forward, they remain constrained by resolution limits, computational cost, and model scalability.
@@
A primary bottleneck is the computational complexity of downscaling at high resolution. For instance, Prithvi achieves 12 km resolution over Europe but is restricted to 50–60 km globally due to the quadratic scaling of ViT self-attention [20]. ViTs divide spatial data into patches, treating each patch as a token. As resolution increases, the number of tokens grows, and self-attention computes pairwise interactions among all tokens, resulting in quadratic growth in memory and compute demands. Unlike Natural Language Processing (NLP) models, which operate on one-dimensional text sequences and scale to over one million tokens [21], ViTs handle high-dimensional spatial inputs with complex dependencies across multiple axes, making long-sequence scaling significantly more computationally intensive. As a result, the longest ViT sequence reported to date is limited to 188k tokens [22]. This constraint directly limits the maximum data resolution ViTs can process, as sequence length scales proportionally with spatial resolution.
@@
Another significant challenge is the uncertainty associated with multi-variable downscaling, as translating coarse-resolution data to fine scales is a highly ill-posed inverse problem—one that becomes even more complex when multiple climate variables are involved. Unlike super-resolution tasks in computer vision [23, 24], where Red-Green-Blue channels represent the same physical quantity, climate variables such as temperature, humidity, and wind are governed by distinct yet interrelated physical processes. This heterogeneity increases the difficulty of learning consistent mappings and exacerbates uncertainty in predictions. A common mitigation strategy is to upsample coarse inputs prior to training [18, 19], which can help reduce uncertainty but significantly increases sequence length and, in turn, computational cost due to ViT’s quadratic complexity. Moreover, upsampling introduces artifacts that can propagate through the model, limiting its effectiveness.
@@
One further limitation of existing downscaling foundation models is their restricted model scale. For example, Prithvi [18] is constrained to 1.4 billion parameters, primarily due to the computational difficulty of scaling ViTs for high-dimensional spatiotemporal data. A major advancement in this area is the Oak Ridge Base AI foundation model for Earth System Predictability (ORBIT) [17], which leverages hybrid sharding and orthogonal parallelisms to scale ViTs to 113 billion parameters—five times larger than previous ViTs and more than 100× larger than typical climate models. While ORBIT represents a major milestone in large-scale Earth system modeling, it is specifically designed for temporal forecasting and does not address spatial downscaling. In particular, it does not resolve the ViT long-sequence bottleneck nor mitigate the uncertainty associated with inverse downscaling problems.
@@
Figure 2: Reslim architecture is split into main and residual paths. No upsampling is used for the main path for ViT training, leading to reduced computations. Residual path is used to condition prediction for reduced uncertainty.
@@
To address the limitations of current foundation models in high-resolution downscaling and to extend the ORBIT framework, we introduce ORBIT-2, a scalable and computationally efficient foundation model for climate downscaling. At its core is a novel ViT architecture, Residual Slim ViT (Reslim), specifically designed to bypass the high computational cost associated with traditional upsampling-based approaches. Unlike existing models that upsample inputs to mitigate uncertainty—resulting in quadratic increases in memory and computation—Reslim operates directly on adaptively compressed spatial inputs, significantly reducing sequence length while preserving critical information. It preserves accuracy and reduces uncertainty through a lightweight residual learning architecture, enabling efficient, low-overhead predictions. Additionally, both training and inference are framed as a Bayesian Estimation problem, incorporating a Markov Random Field Total Variation prior to further constrain uncertainty and improve spatial consistency.
@@
Leveraging the above innovations, ORBIT-2 sets a new benchmark for AI-driven climate and weather downscaling through four key breakthroughs:
  • Efficient Reslim Architecture by operating directly on compressed inputs, achieving over 660$\times$ speedup compared to standard ViTs—without compromising accuracy.
@@
  • Longest ViT Sequence Length by scaling ViT sequence lengths to unprecedented levels—up to 4.2 billion tokens for a 9.5M parameter model and 671 million tokens for a 10B model—surpassing the prior state-of-the-art of 188K tokens by several orders of magnitude [22]. This eliminates the long-standing sequence bottleneck, enabling global downscaling at resolutions as fine as 0.9 Kilometer (km).
@@
  • Scalable Large Model Training by training models with up to 10 billion parameters across 32,768 GPUs, achieving 92–98% strong scaling efficiency and sustained throughput of up to 1.8 ExaFLOPS.
@@
  • State-of-the-Art Accuracy achieving $R^{2}$ scores of 0.98 for precipitation and 0.99 for temperature at 7 km resolution over the continental United States, setting a new standard in high-fidelity downscaling performance.
II Background & State of The Art
@@
Figure 1 illustrates the generalized architecture of leading downscaling foundation models, including Prithvi [18] and ClimateLearn [19]. The inputs consist of low-resolution data with multiple atmospheric physical variables, normalized and bias corrected, and each channel of the architecture reads data for a distinct variable. To address downscaling inverse problem uncertainty, current models upsample coarse-resolution inputs, either via interpolation [19] or convolution [18], before training. This upsampling process is crucial, as it provides a higher-resolution baseline for ViT training, mitigating uncertainty from the inherently ill-posed nature of the multi-variable downscaling problem, thereby improving accuracy and uncertainty. Once upsampled, multi-channel inputs are aggregated into a single-channel representation in feature space, a step that can be performed using either cross-attention mechanisms [25] or shallow convolutional layers [18, 19]. This aggregated representation is then trained by the ViT training blocks, consisting of self-attention and feedforward sub-layers. Finally, the trained output is projected back from feature to image space for each individual physical variable.
@@
This approach, however, introduces major challenges. Upsampling coarse-resolution input data before training increases the sequence length, which increases in proportion to the resolution increase, causing a quadratic increase in memory and computations due to ViT’s self-attention mechanism. This severely limits scalability and resolution, leaving the long-sequence bottleneck unresolved. Prithvi, for example, is limited to relatively coarse 50-60 km resolution for global downscaling. To address this, prior work proposed both AI architecture and High Performance Computing solutions.
@@
Architecture solutions. To mitigate this, architectures like Swin Transformer alleviate some of the computational burden by introducing a hierarchical architecture with shifted window attention [26, 27]. Instead of processing the entire image at once, Swin Transformer partitions the image into smaller, non-overlapping local windows, where self-attention is computed independently within each window. To capture global spatial dependencies, features learned from local windows are aggregated into global features through an architecture hierarchy. While this reduces computing complexity, Swin Transformer has fundamental limitations and its layers of architecture hierarchy must scale proportionally with higher resolution, making it unsuitable for foundation models that needs a single model to generalize across diverse datasets with varying resolutions. Additionally, Swin Transformer’s model size grows with the architecture hierarchy, shifting the computational bottleneck from long-sequence processing to large-model scaling. Consequently, Swin Transformer can only scale up to 147K sequence length on standard 3-channel images [27], far below what is needed for high-resolution, multi-variable downscaling.
@@
Other sparse attention architectures, such as MaxViT [28], attempt to mitigate computational cost by sampling self-attention computations. While this reduces complexity, it comes at the expense of accuracy degradation when the sampling ratio is too high, and it does not address the fundamental quadratic complexity long-sequence problem.
@@
Scaling algorithm solutions. Besides architecture innovations, scaling algorithms, such as sequence parallelism [22, 29, 30], has been proposed as an alternative strategy for scaling ViT sequence length. It distributes image patch tokens across GPUs for parallel computing, alleviating memory constraints. However, because self-attention requires each token to interact with all other tokens from every other GPU, sequence parallelism incurs substantial inter-GPU communication overhead and limits its scalability. More critically, it does not resolve the fundamental quadratic complexity, which causes computational costs to grow rapidly with increased downscaling resolution. As a result, current ViT sequence parallelisms are limited to a maximum of 188K token sequence lengths [22], which remain insufficient for high-resolution multi-variable downscaling.
@@
It is also important to note that other commonly used parallelisms—such as Fully Sharded Data Parallelism (FSDP) [31], Tensor [32], pipeline [33, 34, 35] and hybrid sharded parallelisms [17] are all designed to scale model sizes, rather than long sequences of high-resolution and high-dimensional spatial data. Consequently, none of the existing model parallelisms fundamentally overcome the long-sequence bottleneck in ViTs required for high-resolution global downscaling and there is an urgent need to develop computing efficient and massively parallel architecture and scaling algorithm.
III Innovation Realized
III-A Reslim: A Lightweight ViT Architecture for Scalable and Uncertainty-Aware Downscaling
@@
Main ViT Path. Figure 2 illustrates the Reslim architecture. After tokenizing each low-resolution physical variable into feature embeddings, the model proceeds along two architectural paths: the main ViT and residual paths. Crucially, the main path eliminates input upsampling, avoiding the sequence length inflation and the quadratically increased computing cost typical of ViT architectures.
@@
Adaptive Spatial Compression. Our objective is not only to train directly on low-resolution inputs, but also to further reduce token count and computational cost through compression. Reslim achieves this via an adaptive spatial compression technique, inspired by adaptive image patching and mesh refinement methods [36]. After aggregating multi-variable features (purple block in Fig. 2), the model projects the embedding back into image space and recursively partitions it into spatial quadrants using a quad-tree structure. Partitioning continues for any quadrant where the estimated feature density—computed via Canny edge detection—exceeds a predefined threshold, terminating when a minimum patch size is reached or below predefined threshold.
This approach enables finer-grained learning in feature-rich regions through smaller patches, and coarse-grained learning to smoother regions through larger patches, where less detail is needed. Figure 3 illustrates an example image after variable-aggregated features are mapped back to image space. Compared to conventional uniform patching (Fig.3(a)), where each grid represents an image patch token, the adaptive spatial compression method (Fig.3(b)) reduces the number of patch tokens by 7x in this figure example, significantly decreasing sequence length and computing cost. After ViT training blocks, the decompression module reconstructs the high-resolution output from the compressed embeddings.
@@
Figure 3: Comparison with and without adaptive spatial compression. Each yellow grid is an image patch.
@@
where $y$ is the high-resolution ground-truth, $\hat{x}$ is the Reslim prediction, and $D$ is a latitude weighting matrix to account for the decrease in longitudinal spacing toward the poles. $K$ is the number of output variables, and $N$ is the total number of pixels per variable. The neighborhood $C(\hat{x_{k,i}})$ contains all spatial neighbors of pixel $\hat{x_{k,i}}$, which is the $i^{\textit{th}}$ pixel for the $k^{\textit{th}}$ variable. $b_{i,j}$ is a spatial weighting factor inversely proportional to the euclidean distance between each pixel pair in the same neighborhood. In the above formulation, the first term, $\|y-\hat{x}\|_{D}^{2}$, is the Bayesian forward data likelihood term using a latitude-weighted mean squared error. The second term is a total variation spatial prior, promoting local smoothness by penalizing irregularities within local neighborhoods, but also preserving edges and discontinuities. This makes it well suited for downscaling tasks that depend on spatial coherence and structure preservation.
III-B TILES: Tilewise Efficient Sequence Scaling Algorithm
Figure 4: (a) TILES divides inputs and outputs into tiles, and each GPU downscales a tile separately. (b) Overlapped halos are added to each tile to address border artifacts among neighboring tiles.

### v2

@@
Abstract.
Sparse observations and coarse-resolution climate models limit effective regional decision-making, underscoring the need for robust downscaling. However, existing AI methods struggle with generalization across variables and geographies and are constrained by the quadratic complexity of Vision Transformer (ViT) self-attention. We introduce ORBIT-2, a scalable foundation model for global, hyper-resolution climate downscaling. ORBIT-2 incorporates two key innovations: (1) Residual Slim ViT (Reslim), a lightweight architecture with residual learning and Bayesian regularization for efficient, robust prediction; and (2) TILES, a tile-wise sequence scaling algorithm that reduces self-attention complexity from quadratic to linear, enabling long-sequence processing and massive parallelism. ORBIT-2 scales to 10 billion parameters across 65,536 GPUs, achieving up to 4.1 ExaFLOPS sustained throughput and 74–98% strong scaling efficiency. It supports downscaling to 0.9 km global resolution and processes sequences up to 4.2 billion tokens. On 7 km resolution benchmarks, ORBIT-2 achieves high accuracy with $R^{2}$ scores in range of 0.98–0.99 against observation data.
copyright: none
1. Justification For ACM Gordon Bell Prize for Climate Modelling
ORBIT-2 sets new benchmarks in scalability and scientific impact, training ViT-based models with up to 10 billion parameters on 65,536 GPUs with 74–98% strong scaling efficiency and 4.1 ExaFLOPS sustained throughput in BF16 precision. It enables global climate downscaling at 0.9 km resolution and processes sequences up to 4.2 billion tokens—vastly surpassing prior limits. This breakthrough advances understanding of fine-scale extremes and improves climate modeling critical for adaptation, risk mitigation and decision making.
Attributes Contents
Category
Time-to-solution, Scalability, Throughput
Type of method Dense Vision Transformer Model
Results reported on Whole application including I/O
Precision reported BF16 mixed precision
System scale Measured on Full-Scale System
Measurement mechanism
Timer, Static analysis tool,
FLOP counts
2. Problem Overview
@@
Downscaling bridges this gap by translating coarse resolution global climate model into fine-scale outputs (Giorgi, 2019). This process is critical across many sectors, including agriculture (Jones and Thornton, 2013), water resources (Gutmann et al., 2014), infrastructure planning (Girvetz et al., 2013), energy systems (González-Aparicio et al., 2017), and extreme event forecasting (Hewitson and Crane, 1996). Despite its importance, downscaling remains both scientifically and computationally challenging. It requires physically consistent predictions from massive, high-dimensional spatiotemporal data while maintaining accuracy across diverse regions. These challenges highlight the urgent need for scalable, high-fidelity downscaling approaches to enable effective climate services, disaster preparedness, and policy planning.
@@
Traditional downscaling methods are either dynamical (Giorgi and Gutowski Jr, 2015), which uses nested physical models to simulate fine-scale processes but is computationally intensive, requiring days or weeks on supercomputers and limited to regional domains; or statistical (Wilby and Dawson, 2013; Wilby et al., 1998), which is less computationally demanding but often lacks physical fidelity and generalizability. More recently, artificial intelligence (AI) has emerged as a powerful alternative, providing high-resolution predictions at a fraction of the computational cost-even on edge devices (Koldunov et al., 2024; Kumar et al., 2021; Wang et al., 2021). Most existing AI approaches, however, are task-specific deep learning models (Annau et al., 2023; Eyring et al., 2024; Liu et al., 2024; Aich et al., 2024) that map coarse to fine resolution, requiring retraining for each variable, resolution, or region, and often struggling to generalize across diverse, physically distinct climate variables (Wang et al., 2024).
@@
Figure 1. A generalized AI architecture diagram for state-of-the-art downscaling foundation models. Note that upsampling is used for each channel prior to training blocks to reduce downscaling uncertainty.
@@
To overcome the limitations of task-specific models, recent efforts have introduced foundation models such as Prithvi (Schmude et al., 2024) and ClimateLearn (Nguyen et al., 2023b), which employ multi-task Vision Transformer (ViT)-based architectures to support downscaling across variables and geographic regions. While these models represent an important step forward, they remain constrained by resolution limits, computational cost, and model scalability.
@@
A primary bottleneck is the computational complexity of downscaling at high resolution. For instance, Prithvi achieves 12 km resolution over Europe but is restricted to 50–60 km globally due to the quadratic scaling of ViT self-attention (Han et al., 2023). ViTs divide spatial data into patches, treating each patch as a token, and self-attention computes pairwise interactions among all tokens. As resolution increases, the number of token pairs grows quadratically, resulting in quadratic growth in memory and compute demands. Unlike Natural Language Processing (NLP) models, which operate on one-dimensional text sequences and scale to over one million tokens (Jacobs et al., 2023), ViTs handle high-dimensional spatial inputs with complex dependencies across multiple axes, making long-sequence scaling significantly more computationally intensive. As a result, the longest ViT sequence reported to date is limited to 188k tokens (Tsaris et al., 2024), directly constraining the maximum resolution, as sequence length scales proportionally with spatial resolution.
@@
Another major challenge in downscaling is the inherent uncertainty, as mapping coarse-resolution data to fine scales is a highly ill-posed inverse problem: a single coarse input can correspond to many plausible fine-scale solutions, making the mapping non-unique and sensitive to small perturbations. This challenge is further amplified when downscaling multiple climate variables simultaneously. Unlike super-resolution tasks in computer vision (Li et al., 2022; Lu et al., 2022), where Red-Green-Blue channels represent the same physical quantity, climate variables such as temperature, humidity, and wind are governed by distinct yet interrelated physical processes. This heterogeneity increases the difficulty of learning consistent mappings and exacerbates uncertainty in predictions. A common mitigation strategy is to upsample coarse inputs prior to training (Schmude et al., 2024; Nguyen et al., 2023b), which can help reduce uncertainty but significantly increases sequence length and, in turn, computational cost due to ViT’s quadratic complexity. Moreover, upsampling introduces artifacts that can propagate through the model, limiting its effectiveness.
@@
Existing foundation models also face restricted model scale. For example, Prithvi (Schmude et al., 2024) is constrained to 1.4 billion parameters, primarily due to the computational difficulty of scaling ViTs for high-dimensional spatiotemporal data. A major advancement is the Oak Ridge Base AI foundation model for Earth System Predictability (ORBIT) (Wang et al., 2024), which leverages hybrid sharding and orthogonal parallelisms to scale ViTs to 113 billion parameters—five times larger than previous ViTs and more than 100× larger than typical climate models. However, ORBIT is specifically designed for temporal forecasting and does not address spatial downscaling. In particular, it does not resolve the ViT long-sequence bottleneck nor mitigate the uncertainty associated with inverse downscaling problems.
@@
Figure 2. Reslim architecture is split into main and residual paths. No upsampling is used for the main path for ViT training, leading to reduced computations. Residual path is used to condition prediction for reduced uncertainty.
@@
To overcome these limitations, we introduce ORBIT-2, a scalable and efficient foundation model for climate downscaling. Unlike ORBIT, which focuses on temporal weather forecasting, ORBIT-2 addresses a fundamentally different challenge: high-resolution spatial downscaling. At its core, ORBIT-2 features a novel ViT architecture, Residual Slim ViT (Reslim), specifically designed to bypass the high computational cost associated with traditional upsampling-based approaches. Unlike existing models that upsample inputs to mitigate uncertainty—resulting in quadratic increases in memory and computation—Reslim operates directly on adaptively compressed spatial inputs, significantly reducing sequence length while preserving critical information. It preserves accuracy and reduces uncertainty through a lightweight residual learning architecture, enabling efficient, low-overhead predictions. Additionally, both training and inference are framed as a Bayesian Estimation problem, incorporating a Markov Random Field Total Variation prior to further constrain uncertainty and improve spatial consistency.
@@
Leveraging the above innovations, ORBIT-2 sets a new benchmark through four key breakthroughs:
[…]
@@
  • Longest ViT Sequence Length by scaling ViT sequence lengths to unprecedented levels—up to 4.2 billion tokens for a 9.5M parameter model and 671 million tokens for a 10B model—surpassing the prior state-of-the-art of 188K tokens by several orders of magnitude (Tsaris et al., 2024). This eliminates the long-standing sequence bottleneck, enabling global downscaling at resolutions as fine as 0.9 Kilometer (km).
@@
  • Scalable Large Model Training by training models with up to 10 billion parameters across 65,536 GPUs, achieving 74–98% strong scaling efficiency and sustained throughput of up to 4.1 ExaFLOPS in BF16 precision.
@@
  • State-of-the-Art Accuracy achieving $R^{2}$ scores of 0.98 for precipitation and 0.99 for temperature at 7 km resolution over both the continental United States and global modeling, setting a new standard in high-fidelity downscaling.
3. Background & State of The Art
@@
Fig. 1 illustrates the generalized architecture of leading downscaling foundation models, including Prithvi (Schmude et al., 2024) and ClimateLearn (Nguyen et al., 2023b). The inputs consist of low-resolution data with multiple atmospheric physical variables, normalized and bias corrected, and each channel of the architecture reads data for a distinct variable. To address downscaling inverse problem uncertainty, current models upsample coarse-resolution inputs, either via interpolation (Nguyen et al., 2023b) or convolution (Schmude et al., 2024), before training. This upsampling process is crucial, as it provides a higher-resolution baseline for ViT training, mitigating uncertainty from the inherently ill-posed nature of the multi-variable downscaling problem, thereby improving accuracy and uncertainty. Once upsampled, multi-channel inputs are aggregated into a single-channel representation in feature space, a step that can be performed using either cross-attention mechanisms (Nguyen et al., 2023a) or shallow convolutional layers (Schmude et al., 2024; Nguyen et al., 2023b). This aggregated representation is then trained by the ViT training blocks, consisting of self-attention and feedforward sub-layers. Finally, the trained output is projected back from feature to image space for each individual physical variable.
@@
This approach, however, introduces major challenges. Upsampling coarse-resolution input data before training increases the sequence length, which increases in proportion to the resolution increase, causing a quadratic increase in memory and computations due to ViT’s self-attention mechanism. This severely limits scalability and resolution, leaving the long-sequence bottleneck unresolved. Prithvi, for example, is limited to relatively coarse 50-60 km resolution for global downscaling. To address this, prior work proposed both AI architecture and scaling algorithm solutions.
@@
Architecture solutions. To mitigate this, architectures like Swin Transformer alleviate some of the computational burden by introducing a hierarchical architecture with shifted window attention (Liu et al., 2021, 2022). Instead of processing the entire image at once, Swin Transformer partitions the image into smaller, non-overlapping local windows, where self-attention is computed independently within each window. To capture global spatial dependencies, features learned from local windows are aggregated into global features through an architecture hierarchy. While this reduces computing complexity, Swin Transformer has fundamental limitations and its layers of architecture hierarchy must scale proportionally with higher resolution, making it unsuitable for foundation models that needs a single model to generalize across diverse datasets with varying resolutions. Additionally, Swin Transformer’s model size grows with the architecture hierarchy, shifting the computational bottleneck from long-sequence processing to large-model scaling. Consequently, Swin Transformer can only scale up to 147K sequence length on standard 3-channel images (Liu et al., 2022), far below what is needed for high-resolution, multi-variable downscaling.
@@
Other sparse attention architectures, such as MaxViT (Tu et al., 2022), attempt to mitigate computational cost by sampling self-attention computations. While this reduces complexity, it comes at the expense of accuracy degradation when the sampling ratio is too high, and it does not address the fundamental quadratic complexity long-sequence problem.
@@
Scaling algorithm solutions. Besides architecture innovations, scaling algorithms, such as sequence parallelism (Tsaris et al., 2024; Wang et al., 2023; Lyngaas et al., 2024), has been proposed as an alternative strategy for scaling ViT sequence length. It distributes image patch tokens across GPUs for parallel computing, alleviating memory constraints. However, because self-attention requires each token to interact with all other tokens from every other GPU, sequence parallelism incurs substantial inter-GPU communication overhead and limits its scalability. More critically, it does not resolve the fundamental quadratic complexity, which causes computational costs to grow rapidly with increased downscaling resolution. As a result, current ViT sequence parallelisms are limited to a maximum of 188K token sequence lengths (Tsaris et al., 2024), which remain insufficient for high-resolution multi-variable downscaling.
@@
It is also important to note that other commonly used parallelisms—such as Fully Sharded Data Parallelism (FSDP) (Zhao et al., 2023), Tensor (Shoeybi et al., 2020), pipeline (He et al., 2021; Huang et al., 2019; Kim et al., 2020) and hybrid sharded parallelisms (Wang et al., 2024) are all designed to scale model sizes, rather than long sequences of high-resolution and high-dimensional spatial data. Consequently, none of the existing model parallelisms overcome the long-sequence bottleneck in ViTs required for high-resolution global downscaling and there is an urgent need to develop computing efficient and massively parallel architecture and scaling algorithm.
4. Innovation Realized
4.1. Reslim Architecture
@@
Main ViT Path. Fig. 2 illustrates the Reslim architecture. After tokenizing each low-resolution physical variable into feature embeddings, the model proceeds along two architectural paths: the main ViT and residual paths. Crucially, the main path eliminates input upsampling, avoiding the sequence length inflation and the quadratically increased computing cost typical of ViT architectures.
@@
Adaptive Spatial Compression. Our objective is not only to train directly on low-resolution inputs, but also to further reduce token count and computational cost through compression. Reslim achieves this via an adaptive spatial compression technique, inspired by adaptive image patching and mesh refinement methods (Zhang et al., 2024). After aggregating multi-variable features (purple block in Fig. 2), the model projects the embedding back into image space and recursively partitions it into spatial quadrants using a quad-tree structure. Partitioning continues for any quadrant where the estimated feature density—computed via Canny edge detection—exceeds a predefined threshold, terminating when a minimum patch size is reached or below predefined threshold.
This approach enables finer-grained learning in feature-rich regions through smaller patches, and coarse-grained learning to smoother regions through larger patches, where less detail is needed. Fig. 3 illustrates an example image after variable-aggregated features are mapped back to image space. Compared to conventional uniform patching (Fig.3(a)), where each grid represents an image patch token, the adaptive spatial compression method (Fig.3(b)) reduces the number of patch tokens by 7x in this figure example, significantly decreasing sequence length and computing cost. After ViT training blocks, the decompression module reconstructs the high-resolution output from the compressed embeddings.
@@
Figure 3. Comparison with and without adaptive spatial compression. Each yellow grid is an image patch.
@@
[…]
4.2. TILES: Tilewise Efficient Sequence Scaling Algorithm
Figure 4. (a) TILES algorithm partitions inputs and outputs into tiles, each processed independently on a GPU. Yellow/green pixels denote strong local correlations, while green/blue indicate weaker long-range correlations. (b) Gray halos are added around tiles to avoid border artifacts. Boundary regions (example in red) are duplicated, appearing once in a tile’s halo and once in its neighbor’s interior.